OpenStudy (anonymous):

Five times the difference of twice a number and seven is negative five. Find the number.

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

5*(2x-7)=-5

OpenStudy (anonymous):

5x2=10 10x-35=-5 +35 +35 10x=30 ----------- 10x=10 =3 is this correct

OpenStudy (anonymous):

that is what the eqn says !

OpenStudy (anonymous):

so am i right or wrong?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

help me with another one